Ammonia is produced from its elements in a formation reaction. If 7.81 x 1024 molecules of ammonia are produced, how many mole of ammonia was produced?
1
Expert's answer
2020-06-30T07:46:44-0400

6.023×1023 molecules of ammonia = 1 mole ammonia

7.81×1024 molecules of ammonia = (1×7.81×1024) / 6.023×1023 

                                                       = 12.96 mole ammonia (Answer)

12.96 mole of ammonia was produced.